# G-Loc: Tightly-coupled Graph Localization with Prior Topo-metric Information

### v1

@@ Abstract
Localization in already mapped environments is a critical component in many robotics and automotive applications, where previously acquired information can be exploited along with sensor fusion to provide robust and accurate localization estimates. In this work, we offer a new perspective on map-based localization by reusing prior topological and metric information. Thus, we reformulate this long-studied problem to go beyond the mere use of metric maps. Our framework seamlessly integrates LiDAR, inertial and GNSS measurements, and scan-to-map registrations in a sliding window graph fashion, which allows to accommodate the uncertainty of each observation. The modularity of our framework allows it to work with different sensor configurations (e.g., LiDAR resolutions, GNSS denial) and environmental conditions (e.g., map-less regions, large environments). We have conducted different validation experiments, including deployment in a real-world automotive application, demonstrating the accuracy, efficiency, and versatility of our system in online localization.
@@
Graph Localization, Graph Optimization, Topo-metric Map.
@@ I Introduction
Robust and precise (global) localization is essential for autonomous robots and self-driving vehicles. Simultaneous Localization and Mapping (SLAM) solutions [1, 2, 3] have emerged as a response to 3D reconstruction and online localization, growing exponentially in recent years also in the autonomous driving realm [4]. However, there are two situations where these methods suffer from limitations. On the one hand, they assume that the robot is teleoperated, or that the trajectory is known. Active SLAM tackles this by integrating online path planning into the problem [5]. On the other hand, it is desirable for many robotic applications to navigate repeatedly in previously visited (and mapped) areas without the need of further updating the built model of the environment. For example, autonomous vehicles often navigate familiar routes, and robots typically operate in known environments when not engaged in exploratory tasks. Using a previously built environment model in such situations can scale down the problem of SLAM to map-based localization. This work specifically focuses on the latter, aiming to leverage existing prior information to address localization.
Figure 1: Overview of the proposed graph-based localization method using prior topological and metric information from the environment. Blue elements form the active graph and orange elements the reference graph.
At first glance, without the need of updating an environment model, the use of Global Navigation Satellite System (GNSS) in (semi-)urban areas may seem convenient for localization, although its use alone has long been proven to be neither robust nor accurate enough for the required applications [6]. Combining measurements from multiple sensors and finding correspondences with a reference map model is key to robust, reliable and accurate localization in automotive applications. Most of the existing work tackles this task by using filters for measurement integration, and by exploiting dense high-resolution point cloud maps previously generated by SLAM algorithms [7, 8, 9]. However, modern SLAM systems, based on probability and graph theories [1], contain much more information beyond the resulting reconstruction: a rich underlying graph that captures the localization estimates and accommodates the observation uncertainties. This valuable topological information has been overlooked in the literature.
In this work, we present G-Loc, a novel tightly-coupled graph localization approach that exploits prior topo-metric knowledge of the environment. We associate this information with LiDAR (Light Detection And Ranging) measurements and seamlessly integrate it with IMU (Inertial Measurement Unit) and GNSS data in a robust and efficient manner. We aim to encourage the use of all available information generated by modern SLAM systems,going beyond purely metric maps.
@@ II Related Work
Levinson et al. [6] are among the first to exploit a previously-built map for localization. They correlate LiDAR intensity measurements with the prior map (a grid representation containing a Gaussian distribution of intensity in each cell) using a particle filter. They also incorporate GNSS measurements to limit particle dispersion. However, this and other related work [10, 11], reduce the estimation problem to 3 degrees of freedom, assuming that the rest are estimated, e.g., from an IMU. [12] estimates a global 3D pose instead after incorporating altitude into the distribution of the cells.
@@
A different approach is followed in [13], where the reference map consists of road primitives. An unscented Kalman filter is used to combine GNSS, odometry and the matching between the prior map and the online primitives extracted from images via convolutional neural networks. The challenge of maintaining and matching a dense point cloud map is also addressed in [14], where such representations are transformed into a sparse multi-layer vector map, and [15], where features are extracted from lateral galleries in a tunnel-like environment.
@@
More recently, advances in point cloud registration have brought these methods to the forefront of map-based localization, whether feature-based or not. Yoneda and Mita [7] first matched LiDAR measurements to a geo-referenced point cloud map using Iterative Closest Point (ICP) [16]. The limitations of this single sensor approach and the sensitivity of ICP to the initial guess are addressed in subsequent works, all of them being filter-based [8, 17, 9, 18, 19, 20]. Nagy and Benedek [9] fuse point cloud and semantic matching with GNSS measurements to localize a robot in a high resolution 3D semantic cloud. In [21], instead of matching the online LiDAR data to the reference point cloud map, the latter is rendered in 2D; this allows to match monocular images. Autoware [8], a widely used open-source library constantly under development, features global localization in a prior map (either in the form of a point cloud or a high-definition vector map). This approach relies on aligning the input clouds with the reference map using the Normal Distribution Transform (NDT) [22] technique. The optimized transformation is integrated with IMU and GNSS measurements within an extended Kalman filter module. PoseMap [23] incorporates the trajectory into the prior model, thus going beyond purely metric representations. Each of the robot poses is assigned a submap of surfels and the alignment of the online cloud is done only with the $k$-closest neighbors, significantly reducing the computational cost and allowing to take advantage of the full LiDAR range.
@@
In contrast to the filter-based techniques mentioned above, a handful of methods have recently started to use a graph formulation, borrowing from the advances in graph SLAM and due to its improved robustness and accuracy [24]. Dubé et al. [25] estimate the robot’s pose in a prior global map by extracting segments from a 3D point cloud and matching neighbors in the feature space. The resulting transformations are later fed into a pose-graph SLAM framework along with pre-integrated IMU measurements. [26] also employs a pose-graph formulation for localization, combining wheel odometry, GNSS, and semantic matching with the reference map.
In parallel to our work, Koide et al. [27] recently proposed a map-based localization approach that tightly couples range and IMU data within a sliding window factor graph optimization.
The localization method proposed in this letter falls into the graph-based category. In contrast to previous work, we combine GNSS data, IMU pre-integrated measurements, LiDAR odometry and map-matching constraints within a graph optimization framework, thus achieving global localization in a robust and efficient manner. Furthermore, we go beyond dense metric representations (unlike [25, 26, 27]) and fully exploit the prior knowledge obtained during the mapping task, somewhat following [23] but leveraging the pose-graph topology instead of just the trajectory poses. This allows us not only to relax the neighborhood search and restrict the alignment of the online clouds to certain regions of the reference map (i.e., submaps), but also to integrate those neighboring reference nodes into the optimization framework (see Figure 1). G-Loc is a modular and optimized framework that adapts to a wide variety of algorithms and sensors with minimal parameter tuning and is capable of handling large-scale environments. We have validated the proposed method on several datasets, as well as in a real automotive application, demonstrating the robustness and accuracy of G-Loc even in map-less regions compared to other popular state-of-the-art methods.
@@ III Problem Formulation
First of all, let us denote the map and robot base frames by $m$ and $r$, respectively. To simplify the notation, all the sensors will be assumed to be in $r$. In its most basic form, map-based 3D localization can be formulated as the problem of estimating the robot poses that best match the sensor observations with each other and with the map, i.e., finding the maximum a posteriori (MAP) estimates
@@
where $p(\cdot)$ denotes probability, $\mathcal{M}$ is the prior map, $\mathcal{Z}$ is the set of sensor measurements, and $\mathcal{T}\triangleq\{\boldsymbol{}^{m}{T}_{r_{i}}\in SE(3)\,|\,i\in\mathcal{K}_{t}\}$ is the set of relative transformations between the map and the latest $N_{\mathcal{K}}\texttt{+}1$ robot frames. To maintain tractability, the set of frames $\mathcal{K}_{t}\triangleq\{\max(0,t\texttt{-}N_{\mathcal{K}}),\dots,t\}$ has been limited in a sliding window fashion.
On the one hand, we assume that $\mathcal{M}$ has been previously built by a graph-SLAM algorithm, and thus contains metric and topological information (in the form of point clouds and a pose-graph, respectively). Furthermore, the metric part is associated with the nodes of the graph in such a way that each vertex is assigned a set of point clouds (or submap, $\mathcal{S}$) of its surroundings (see upper part of Figure 1). Then:
@@
$\mathcal{S}_{\ell}\triangleq\{\mathcal{L}_{q}^{\ell}\,|\,q=1,\dots,N_{\mathcal{S}_{\ell}}\}\,,$ (2c)
where $\mathcal{R}\triangleq\{0,\dots,\dim(\mathcal{V})\}$ is the reference frame set, $(\mathcal{V},\mathcal{E})$ are the sets of vertices and edges in the pose-graph $\mathcal{G}$, and any given point cloud $\mathcal{L}\triangleq\{\tilde{\boldsymbol{p}}_{l}\in\mathbb{R}^{3}\,|\,l\,\texttt{=}\,1,\dots,N_{\mathcal{L}}\}$. Notice that the submaps of nearby vertices might overlap and that elements of the pose-graph will live in $SE(3)$. Additionally, the graph may contain GNSS information, making also the map geo-referenced and aligned to the East, North, Up (ENU) global coordinate system.
On the other hand, the set of measurements collected up to time $t$ consists of LiDAR point clouds, IMU observations between consecutive frames, and GNSS data:
$\mathcal{Z}_{t}\triangleq\{\mathcal{Z}^{\text{LiDAR}}_{i},\mathcal{Z}^{\text{IMU}}_{i,j},\mathcal{Z}^{\text{GNSS}}_{i}\}_{i,j\in\mathcal{K}_{t}}\,.$ (3)
@@
where $\|\boldsymbol{e}\|^{2}\triangleq(\boldsymbol{e}^{T}\boldsymbol{\Sigma}^{\texttt{-}1}\boldsymbol{e})\in\mathbb{R}$ is the quadratic error for a measurement with covariance $\boldsymbol{\Sigma}$. $\boldsymbol{e}_{i,j}^{\text{LIO}}$ are the residual errors associated with the LiDAR-inertial odometry (LIO) estimates (described in Section IV-A), $\boldsymbol{e}_{i}^{\mathcal{M}}$ are those associated with the correspondences between LiDAR observations and the prior map (Section IV-B), and $\boldsymbol{e}_{i}^{\text{GNSS}}$ are the translation errors associated with the GNSS measurements.
@@ IV Methods
G-Loc is divided into three main threads, namely: LIO, Localization and Dynamic Loading, to which the following subsections are dedicated. Figure 2 contains an overview of the system, illustrating the different modules and their connections. Thanks to its parallelized architecture, the entire system runs in real time within a ROS2 Humble [29] framework. It is able to use GPU-accelerated cloud registration methods, and it is robust to a wide variety of sensor configurations with minimal parameter tuning.
@@ IV-A LiDAR-inertial Odometry (LIO)
The de-skewed cloud is then aligned with the previous submap $\mathcal{S}_{t-1}$ created by concatenating the $N_{\mathcal{S}}$ previous point clouds (see Figure 3). Scan-to-submap alignment relies either on ICP or NDT techniques. Our system supports a variety of registration methods, but two have proven to be superior: (i) the efficient and GPU-accelerated implementation of voxelized generalized ICP [31], and (ii) the parallelized version of NDT [^1] https://github.com/koide3/ndt_omp. This process outputs LiDAR odometry (LO) estimates between frames (i.e., the transformations ${}^{i\texttt{-}1}\tilde{\boldsymbol{T}}^{\text{LO}}_{i}$), which can be introduced into the graph as prior edges to constrain the pose of the inertial state after computing ${}^{i}\tilde{\boldsymbol{T}}^{\text{LO}}=\left(\boldsymbol{T}_{i\texttt{-}1}\ {}^{i\texttt{-}1}\tilde{\boldsymbol{T}}^{\text{LO}}_{i}\right)^{\texttt{-}1}$. The residuals for the $i$-th frame will be given by:
@@
Simultaneously with the above, IMU measurements between LiDAR frames are pre-integrated following [32]. Pre-integration provides an estimate of the robot’s motion between successive frames, also allowing the scan-to-submap matching to be fed with a decent initial guess (see Figure 3). IMU data is efficiently incorporated into the graph optimization framework by creating an inertial edge [30, 2] between frames; this edge is actually composed of several different edges that correlate the poses and velocities between frames, and also model the bias variation over time (assumed to follow a Brownian motion pattern). Acceleration and rotation rate measurements ($\tilde{\boldsymbol{a}}$ and $\tilde{\boldsymbol{\omega}}$) can be integrated between $i$-th and $j$-th LiDAR frames to estimate the evolution of the sensor’s pose and velocity. Let us consider the motion model:
$\boldsymbol{R}_{j}$ $=\boldsymbol{R}_{i}\textstyle\prod\limits_{k=i}^{j-1}\exp((\tilde{\boldsymbol{\omega}}_{k}\texttt{-}\boldsymbol{b}_{k}^{g}\texttt{-}\boldsymbol{\eta}_{k}^{g})\Delta t)\,,$
$\boldsymbol{v}_{j}$ $=\boldsymbol{v}_{i}+\boldsymbol{g}\Delta t+\textstyle\sum\limits_{k=1}^{j-1}\boldsymbol{R}_{k}(\tilde{\boldsymbol{a}}_{k}\texttt{-}\boldsymbol{b}_{k}^{a}\texttt{-}\boldsymbol{\eta}_{k}^{a})\Delta t\,,$
$\boldsymbol{p}_{j}$ $=\boldsymbol{p}_{i}+\textstyle\sum\limits_{k=1}^{j-1}\left(\boldsymbol{v}_{k}\Delta t+\frac{1}{2}\boldsymbol{g}\Delta t^{2}+\frac{1}{2}\boldsymbol{R}_{k}(\tilde{\boldsymbol{a}}_{k}\texttt{-}\boldsymbol{b}_{k}^{a}\texttt{-}\boldsymbol{\eta}_{k}^{a})\Delta t^{2}\right)\,,$ (7)
with $\Delta t$ the time between the two frames, $\boldsymbol{\eta}^{g},\boldsymbol{\eta}^{a}\in\mathbb{R}^{3}$ additive white noise and $\boldsymbol{T}\triangleq(\boldsymbol{R},\boldsymbol{p})$, where $\boldsymbol{R}\in SO(3)$ and $\boldsymbol{p}\in\mathbb{R}^{3}$. From (7), we can define the relative increments between frames and the preintegrated measurement model, i.e., $\Delta\tilde{\boldsymbol{R}}_{ij}$, $\Delta\tilde{\boldsymbol{v}}_{ij}$ and $\Delta\tilde{\boldsymbol{p}}_{ij}$ (see [32] for details). Using them, the residual error of the IMU measurements $\boldsymbol{e}^{\text{IMU}}\triangleq(\boldsymbol{e}_{\Delta\boldsymbol{R}},\boldsymbol{e}_{\Delta\boldsymbol{v}},\boldsymbol{e}_{\Delta\boldsymbol{p}},\boldsymbol{e}_{\Delta\boldsymbol{b}})^{T}\in\mathbb{R}^{15}$ will be given by [32]:
@@
$\mathcal{X}^{\star}=\underset{\mathcal{X}}{\mathrm{argmin}}\left(\sum_{i\in\mathcal{I}_{t}}\|\boldsymbol{e}_{i}^{\text{LO}}\|^{2}+\sum_{i,j\in\mathcal{I}_{t}}\|\boldsymbol{e}_{i,j}^{\text{IMU}}\|^{2}\right)$ (9)
High-frequency odometry (at IMU rate) can be obtained by composing the LO with the relative motion predictions from the preintegrated measurement model.
The success of any registration algorithm depends on a good initial guess and sufficient overlap between the input clouds. The first issue has already been addressed, and the second is mitigated by the submaps, which in turn require some maintenance. After adding a new point cloud to the submap, it is voxelized (e.g., via uniform sampling) to prune redundant points. To restrain complexity, the submap size is limited to include the last $N_{\mathcal{S}}$ point clouds. A side effect of this approach is high adaptability to different range sensors: we can handle densities from 16 to 128 beams and different patterns by changing $N_{\mathcal{S}}$ (e.g., submaps containing one single cloud are suitable for 128-plane LiDARs). This provides the flexibility to work with different sensors with minimal parameter tuning. By effectively configuring the submaps, we can also keep the computational cost within reasonable limits for the resources available and the precision required in each application.
Figure 4: Visualization of G-Loc. The prior graph is shown in orange and the active graph in blue (vertices are represented as squares and edges as lines connecting them). The upper left corner displays a zoom of the active graph, which consists of the sliding graph and a subgraph of the prior model. The online LiDAR cloud is shown as red points, while the loaded submaps of the prior model are colored by height. Note that only a subset of the metric model (i.e., the submaps corresponding to vertices within the robot’s neighborhood) has been loaded.
@@
The main (global) localization process is based on optimizing a sliding graph together with regions of interest of the reference graph (see Figure 1).
The localization process starts with estimating the robot’s initial pose within the prior map using GNSS measurements [^2] In the absence of GNSS, this initial estimate can be given manually by the user, e.g., using a ROS2 service.. It then queries the prior topo-metric model to find the $k$-closest reference vertices (for the exemplary graph shown in Figure 1, at time $t$, it will be $\boldsymbol{T}^{\prime}_{\ell}$). Next, the current 3D LiDAR scan is aligned with the submaps associated with these vertices ($\mathcal{S}_{\ell}^{\prime}$). The active graph combines the constraints enforced by the results of these alignments along with the relative motion data provided by the LIO module and GNSS measurements (if available). Therefore, this new graph will contain vertices from both the prior reference graph and the online sliding graph; the latter encodes current and past robot states in a sliding window approach (i.e., $\mathcal{T}$). The size of the sliding graph is limited to $\dim(\mathcal{T})$), being older nodes marginalized as they leave the window. Figure 4 displays the active (blue) and reference (orange) graphs on top of the prior point cloud map for an example localization task. The online cloud is shown in red. As shown in the zoomed area, multiple reference vertices can be associated with a single vertex in the sliding graph.
The optimization of the active graph is governed by (4), subject to:
@@
$\boldsymbol{e}_{i}^{\mathcal{M}}$ $\triangleq\sum\limits_{k\in\mathcal{R}_{i}}\log\left(\boldsymbol{T}_{i}\ {}^{i}\tilde{\boldsymbol{T}}_{k}^{\mathcal{M}}\ (\boldsymbol{T}^{\prime}_{k})^{\texttt{-}1}\right)^{\vee}\in\mathbb{R}^{6}\,,$ (10b)

### v2

@@ Abstract
Localization in already mapped environments is a critical component in many robotics and automotive applications, where previously acquired information can be exploited along with sensor fusion to provide robust and accurate localization estimates. In this work, we offer a new perspective on map-based localization by reusing prior topological and metric information. Thus, we reformulate this long-studied problem to go beyond the mere use of metric maps. Our framework seamlessly integrates LiDAR, inertial and GNSS measurements, and cloud-to-map registrations in a sliding window graph fashion, which allows to accommodate the uncertainty of each observation. The modularity of our framework allows it to work with different sensor configurations (e.g., LiDAR resolutions, GNSS denial) and environmental conditions (e.g., mapless regions, large environments). We have conducted several validation experiments, including the deployment in a real-world automotive application, demonstrating the accuracy, efficiency, and versatility of our system in online localization.
@@
Robot Localization, Graph Optimization, Topo-metric Map.
@@ I Introduction
Robust and precise (global) localization is essential for autonomous robots and self-driving vehicles. Simultaneous Localization and Mapping (SLAM) solutions [1, 2] have emerged as a response to 3D reconstruction and online localization, growing exponentially in recent years also in the autonomous driving realm [3]. However, there are two situations where these methods suffer from limitations. On the one hand, they assume that the robot is teleoperated, or that the motion actions are known a priori. Active SLAM tackles this by integrating online path planning into the problem [4]. On the other hand, it is desirable for many robotic applications to navigate repeatedly in previously visited (and mapped) areas without the need of further updating the built model of the environment. For example, autonomous vehicles often navigate familiar routes, and robots typically operate in known environments when not engaged in exploratory tasks. Using a previously built environment model in such situations can scale down the problem of SLAM to map-based localization. This work specifically focuses on the latter, aiming to leverage existing prior information to address localization.
Figure 1: Visualization of G-Loc in a sequence of the Newer College Extension dataset [5]. The prior topological (orange pose-graph) and metric (colored pointcloud) model is exploited for online robust localization. An active graph (blue) is optimized, containing certain vertices from the prior graph and the most recent robot states.
At first glance, without the need of updating an environment model, the use of Global Navigation Satellite System (GNSS) in (semi-)urban areas may seem convenient for localization, although its use alone has long been proven to be neither robust nor accurate enough for the required applications [6]. Combining measurements from multiple sensors and finding correspondences with a reference map model is key to robust, reliable and accurate localization in automotive applications. Most of the existing work tackles this task by using filters for measurement integration, and by exploiting dense high-resolution point cloud maps previously generated by SLAM algorithms [7, 8, 9, 10]. However, modern SLAM systems, based on probability and graph theories [1], contain much more information beyond the resulting reconstruction: a rich underlying graph that captures the localization estimates and accommodates the observation uncertainties. This valuable topological information has been overlooked in the literature.
In this work, we present G-Loc (see Figure 1), a novel tightly-coupled graph localization approach that exploits prior topo-metric knowledge of the environment. We propose to go beyond purely geometric maps and use all the information provided by modern SLAM systems. In addition, we efficiently fuse this information with LiDAR (Light Detection And Ranging), IMU (Inertial Measurement Unit), and GNSS observations to create a robust and accurate localization system.
@@ II Related Work
Levinson et al. [6] are among the first to exploit a previously-built map for localization. They correlate LiDAR intensity measurements with the prior map (a grid representation containing a Gaussian distribution of intensity in each cell) using a particle filter. They also incorporate GNSS measurements to limit particle dispersion. However, this and other related work [11], reduce the estimation problem to 3 degrees of freedom, assuming that the rest are estimated, e.g., from an IMU. [12] estimates a global 3D pose instead after incorporating altitude into the distribution of the cells.
@@
A different approach is followed in [13], where the reference map consists of road primitives. An Unscented Kalman Filter (UKF) is used to combine GNSS, odometry and the matching between the prior map and the online primitives extracted from images via convolutional neural networks. The challenge of maintaining and matching a dense point cloud map is also addressed in [14], where such representations are transformed into a sparse multi-layer vector map, and [15], where features are extracted from lateral galleries in tunnel-like environments.
@@
More recently, advances in point cloud registration have brought these methods to the forefront of map-based localization, whether feature-based or not. Yoneda and Mita [7] first matched LiDAR measurements to a georeferenced point cloud map using Iterative Closest Point (ICP) [16]. The limitations of this single sensor approach and the sensitivity of ICP to the initial guess are addressed in subsequent works, all of them being filter-based [8, 9, 17, 18, 10]. Nagy and Benedek [9] fuse point cloud and semantic matching with GNSS measurements to localize a robot in a high resolution 3D semantic cloud. In [19], instead of matching the online LiDAR data to the reference point cloud map, the latter is rendered in 2D; this allows to match monocular images. In [20] LiDAR Odometry (LO) and IMU data are combined in a UKF to improve the matching of the current scan with a portion of the prior point cloud map; this portion is given by cropping the map around the GNSS pose estimation. Autoware [8], a widely used open-source library constantly under development, features global localization in a prior map (either in the form of a point cloud or a high-definition vector map). This approach relies on aligning the input clouds with the reference map using the Normal Distribution Transform (NDT) [21] technique. The optimized transformation is integrated with IMU and GNSS measurements within an extended Kalman filter module. Xia et al. [22] combine GNSS, IMU and NDT-based map matching on a Kalman filter. However, they use a light map instead of selecting regions of interest from a dense map. On the downside, this requires a special offline process. PoseMap [23] incorporates the trajectory into the prior model. Each robot pose is assigned a submap of surfels and the online cloud is aligned only with the $k$-closest neighbors, reducing the computational cost and allowing to use the full LiDAR range.
@@
In contrast to the filter-based techniques mentioned above, a handful of methods have recently started to use a graph formulation, borrowing from the advances in graph SLAM and due to its improved robustness and accuracy [24]. Dubé et al. [25] estimate the robot’s pose in a prior global map by extracting segments from a 3D point cloud and matching neighbors in the feature space. The resulting transformations are later fed into a pose-graph SLAM framework along with pre-integrated IMU measurements. [26] also employs a pose-graph formulation for localization, combining wheel odometry, GNSS, and semantic matching with the reference map. In parallel to our work, Koide et al. [27] recently proposed a map-based localization approach that tightly couples range and IMU data within a sliding window factor graph optimization.
The localization method proposed in this letter falls into the graph-based category. G-Loc is a modular and optimized framework that adapts to a wide variety of sensor and environment configurations, capable of handling large-scale environments, GNSS denial, mapless regions, etc. Unlike previous work [25, 26, 27] we go beyond geometric representations and fully exploit the knowledge gathered during a previous mapping task, somewhat following [23] but leveraging the pose-graph topology instead of just the trajectory poses. More specifically, our main contributions are as follows:
  • A novel approach to map-based localization that exploits the topology of the prior model both for an efficient search of regions of interest, and for accommodating their uncertainty in a graph optimization process.
  • A complete localization system that combines precise LiDAR-inertial odometry, GNSS and map matching constraints. It can accurately locate a robot under various sensor configurations and environmental challenges.
  • Thorough experiments demonstrating the accuracy and versatility of our system compared to state-of-the-art localization and SLAM methods, as well as real world deployment in an autonomous driving scenario.
@@ III Problem Formulation
First of all, let us denote the map and robot base frames by $m$ and $r$, respectively. To simplify the notation, all the sensors will be assumed to be in $r$. In its most basic form, map-based 3D localization can be formulated as the problem of estimating the robot poses that best match the sensor observations with each other and with the map, i.e., finding the maximum a posteriori estimates
@@
where $p(\cdot)$ denotes probability, $\mathcal{M}$ is the prior map, $\mathcal{Z}$ is the set of sensor measurements, and $\mathcal{T}\triangleq\{{}^{m}\boldsymbol{T}_{r_{i}}\in SE(3)\,|\,i\in\mathcal{K}_{t}\}$ is the set of relative transformations between the map and the latest $N_{\mathcal{K}}\texttt{+}1$ robot frames. To maintain tractability, the set of frame indices $\mathcal{K}_{t}\triangleq\{\max(0,t\texttt{-}N_{\mathcal{K}}),\dots,t\}$ has been limited in a sliding window fashion.
On the one hand, we assume that $\mathcal{M}$ has been previously built by a graph-SLAM algorithm, and thus contains metric and topological information (in the form of point clouds and a pose-graph, respectively). Furthermore, the metric part is associated with the nodes of the graph in such a way that each vertex is assigned a local map (or submap, $\mathcal{S}^{\prime}$) of its surroundings (see upper part of Figure 4). Then:
@@
where $\mathcal{R}\triangleq\{0,\dots,\dim(\mathcal{V})\}$ is the reference frame index set, and $(\mathcal{V},\mathcal{E})$ are the sets of vertices and edges in the pose-graph. The $\ell$-th submap will be formed by concatenating the point clouds between the $\ell$ and $\ell-N_{s}$ LiDAR frames, as described in [29], or just a region around the reference vertex. The latter can be easily obtained by splitting the resulting map of a typical SLAM process [30] into grids. Notice that the submaps of nearby vertices might overlap and that elements of the pose-graph will live in $SE(3)$. Additionally, the graph may contain GNSS information, making also the map georeferenced and aligned to the East, North, Up (ENU) global coordinate system. On the other hand, the set of measurements collected up to time $t$ consists of LiDAR point clouds, IMU observations between consecutive frames, and GNSS data: $\mathcal{Z}_{t}\triangleq\{\mathcal{Z}^{\text{LiDAR}}_{i},\mathcal{Z}^{\text{IMU}}_{i,j},\mathcal{Z}^{\text{GNSS}}_{i}\}_{i,j\in\mathcal{K}_{t}}$.
@@
where $\|\boldsymbol{e}\|^{2}\triangleq(\boldsymbol{e}^{T}\boldsymbol{\Sigma}^{\texttt{-}1}\boldsymbol{e})\in\mathbb{R}$ is the quadratic error for a measurement with covariance $\boldsymbol{\Sigma}$. $\boldsymbol{e}_{i,j}^{\text{LIO}}$ are the residual errors associated with the LiDAR-inertial odometry (LIO) estimates, $\boldsymbol{e}_{i}^{\mathcal{M}}$ are those associated with the correspondences between LiDAR observations and the prior map, and $\boldsymbol{e}_{i}^{\text{GNSS}}$ are the translation errors associated with the GNSS measurements.
@@ IV Methods
G-Loc is divided into three main threads: LIO, Localization and Dynamic Loading, to which the following subsections are dedicated. Figure 2 provides an overview of the system, showing the different modules and their connections. Thanks to its parallelized architecture, the entire system runs in real time within a ROS2 Humble [31] framework. It is able to use GPU-accelerated cloud registration and is robust to a wide range of sensor configurations with minimal parameter tuning.
@@ IV-A LiDAR-inertial Odometry (LIO)
The de-skewed cloud is then aligned with the previous submap $\mathcal{S}_{t-1}$ created by concatenating the $N_{\mathcal{S}}$ previous point clouds (see Figure 3). Scan-to-submap alignment relies either on the efficient and GPU-accelerated implementation of voxelized generalized ICP (FastGICP) [32], or the parallelized version of NDT (NDTOmp) [33]. This process outputs LO estimates between frames (i.e., the transformations ${}^{i\texttt{-}1}\tilde{\boldsymbol{T}}^{\text{LO}}_{i}$), which can be introduced into the graph as prior edges to constrain the pose of the inertial state after computing ${}^{i}\tilde{\boldsymbol{T}}^{\text{LO}}=\left(\boldsymbol{T}_{i\texttt{-}1}\ {}^{i\texttt{-}1}\tilde{\boldsymbol{T}}^{\text{LO}}_{i}\right)^{\texttt{-}1}$. The residuals for the $i$-th frame will be given by:
@@
Simultaneously with the above, IMU measurements between LiDAR frames are pre-integrated following [34]. Pre-integration provides an estimate of the robot’s motion between successive frames, also allowing the scan-to-submap matching to be fed with a decent initial guess (see Figure 3). High-frequency odometry (at IMU rate) can be obtained by composing the LO with the relative motion predictions from the preintegrated measurement model. IMU data is efficiently incorporated into the graph optimization framework by creating an inertial edge [30, 2] between frames; this edge is actually composed of several different edges that correlate the poses and velocities between frames, and also model the bias variation over time (assumed to follow a Brownian motion pattern). We use the method proposed in [34] to define the relative increments between frames and the preintegrated measurement model, i.e., $\Delta\tilde{\boldsymbol{R}}_{ij}$, $\Delta\tilde{\boldsymbol{v}}_{ij}$ and $\Delta\tilde{\boldsymbol{p}}_{ij}$ (see [34] for details). This allows us to define their residual error as $\boldsymbol{e}^{\text{IMU}}\triangleq(\boldsymbol{e}_{\Delta\boldsymbol{R}},\boldsymbol{e}_{\Delta\boldsymbol{v}},\boldsymbol{e}_{\Delta\boldsymbol{p}},\boldsymbol{e}_{\Delta\boldsymbol{b}})^{T}\in\mathbb{R}^{15}$, where:
@@
$\mathcal{X}^{\star}=\underset{\mathcal{X}}{\mathrm{argmin}}\left(\textstyle\sum\limits_{i\in\mathcal{I}_{t}}\|\boldsymbol{e}_{i}^{\text{LO}}\|^{2}+\sum\limits_{i,j\in\mathcal{I}_{t}}\|\boldsymbol{e}_{i,j}^{\text{IMU}}\|^{2}\right)$ (7)
The success of registration algorithms depend on a good initial guess and sufficient overlap between the input clouds. The first issue has already been addressed, and the second is mitigated by the submaps, which in turn require some attention. We use the scan-to-submap alignment to add new point clouds to the submap. Then, the latter is downsampled (e.g., via uniform sampling) to prune redundant points. To restrain complexity, the submap size is limited to include the last $N_{\mathcal{S}}$ point clouds. A side effect of this approach is high adaptability to different range sensors: we can handle densities from 16 to 128 beams and different patterns by changing $N_{\mathcal{S}}$ By effectively configuring the submaps (e.g., $N_{\mathcal{S}}=1$ for 128-plane LiDARs), we keep the computational cost manageable while achieving the precision required for each application.
@@
The global localization process is based on optimizing the active graph, which composed of a sliding graph with the latest robot states ($\boldsymbol{T}$), and regions of interest of the reference graph (i.e., certain reference vertices, $\boldsymbol{T}^{\prime}$). See Figure 4.
Figure 4: Overview of the proposed graph-based localization method using prior topological and metric information. Orange elements form the reference graph, and blue elements form the active graph.
The localization process starts with estimating the robot’s initial pose within the prior map using GNSS measurements [^1] In the absence of GNSS, this initial estimate can be given manually by the user, e.g., using a ROS2 service.. It then queries the prior topo-metric model to find the $k$-closest reference vertices (for the exemplary graph shown in Figure 4, at time $t$, they will be $\boldsymbol{T}^{\prime}_{\ell}$ and $\boldsymbol{T}^{\prime}_{\ell+1}$). Next, the current scan is aligned with the submaps associated with these vertices ($\mathcal{S}_{\ell}^{\prime},\mathcal{S}_{\ell+1}^{\prime}$). The active graph combines the constraints enforced by the results of these alignments with the relative motion data provided by the LIO and GNSS measurements (if available). Therefore, this new graph will contain vertices from both the prior reference graph and the online sliding graph. The latter encodes current and past robot states in a sliding window approach (i.e., $\mathcal{T}$), being older nodes marginalized as they leave the window. Figure 1 displays the active (blue) and reference (orange) graphs on top of the prior point cloud map for a localization task. The online cloud is shown in red. Note how multiple reference vertices can be associated with a single vertex in the sliding graph (as is also shown in Figure 4).
The optimization of the active graph is governed by (3), s.t.:
@@
$\boldsymbol{e}_{i}^{\mathcal{M}}$ $\triangleq\textstyle\sum\limits_{k\in\mathcal{R}_{i}}\log\left(\boldsymbol{T}_{i}\ {}^{i}\tilde{\boldsymbol{T}}_{k}^{\mathcal{M}}\ (\boldsymbol{T}^{\prime}_{k})^{\texttt{-}1}\right)^{\vee}\in\mathbb{R}^{6}\,,$ (8b)